A long narrow horizontal slit is placed 1 mm above a horizontal plane mirror. The interference between the light coming directly from the slit and that after reflection is seen on a screen 1.0 m away from the slit. Find the fringe-width if the light used has a wavelength of 700 nm.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Given that, D = 1 m, λ = 700 nm = 700 * 10^–9 m Since, a = 2 mm, d = 2a = 2mm = 2 * 10^–3 m (L loyd’s mirror experiment) Fringe width = λD/d = 700 * 10^-9m * 1m/2 * 10^-3m= 0.35 mm.